OpenStudy (anonymous):

-3x-2y=5.5; x+3y=7.5

OpenStudy (anonymous):

What are you stuck on?

OpenStudy (anonymous):

i know you solve it by substitution, i cant figure out which equation to start off with

OpenStudy (anonymous):

-3x-2y=5.5 (x+3y=7.5) *3

OpenStudy (anonymous):

-3x-2y=5.5 + 3x+9y=22.5 The x's cancel, so you get 7y=28

OpenStudy (anonymous):

solve for y, then input the value of y into (x+3y)=7.5 to solve for x

OpenStudy (anonymous):

@smarks01